# Efficient Policy Optimization in Robust Constrained MDPs with Iteration Complexity Guarantees

### v2

@@ Appendix A Proof of Proposition 1
$\Phi(\pi)-\Phi(\pi^{*})\leq J_{c_{i}}^{\pi,p}-J_{c_{i}}^{\pi^{*}}\leq J_{c_{i}}^{\pi,p}-J_{c_{i}}^{\pi^{*}}.$ (30)
We now bound the right-hand side.
$J_{c_{i}}^{\pi,p}-J_{c_{i}}^{\pi^{*},p}=V^{\pi,p}_{c_{i}}(\rho)-\mathbb{E}_{s\sim\rho}\mathbb{E}_{\pi^{*}}[r(s,a)+\gamma\sum_{s^{\prime}}p(s^{\prime}|s,a)V^{\pi^{*},p}_{c_{i}}(s^{\prime})]$
$=\mathbb{E}_{s\sim\rho}\mathbb{E}_{a\sim\pi^{*}}[V_{c_{i}}^{\pi,p}(\rho)-r(s,a)+\gamma\sum_{s^{\prime}}p(s^{\prime}|s,a)V^{\pi,p}_{c_{i}}(s^{\prime})]$
$-\mathbb{E}_{s\sim\rho}\mathbb{E}_{\pi^{*}}\gamma[\sum_{s^{\prime}}p(s^{\prime}|s,a)V^{\pi^{*},p}_{c_{i}}(s^{\prime})-\sum_{s^{\prime}}p(s^{\prime}|s,a)V^{\pi,p}_{c_{i}}(s^{\prime})]$
$=\sum_{s\sim\rho}\langle Q_{c_{i}}^{\pi},\pi-\pi^{*}\rangle-\gamma\mathbb{E}_{s\sim\rho}\mathbb{E}_{\pi^{*}}[\sum_{s^{\prime}}p(s^{\prime}|s,a)(V^{\pi^{*},p}_{c_{i}}(s^{\prime})-V^{\pi,p}_{c_{i}}(s^{\prime}))]$
$\leq\sum_{s\sim\rho}\langle Q_{c_{i}}^{\pi},\pi-\pi^{*}\rangle+\beta\mathbb{E}_{s\sim\rho}\mathbb{E}_{\pi^{*}}[\sum_{s^{\prime}}p_{0}(s^{\prime}|s,a)(V^{\pi}_{c_{i}}(s^{\prime})-V^{\pi^{*}}_{c_{i}}(s^{\prime}))].$ (31)
where we use the fact that $V_{c_{i}}^{\pi,p}(s^{\prime})\geq V_{c_{i}}^{\pi^{*},p}(s^{\prime})$ as $J_{c_{i}}^{\pi,p}-J_{c_{i}}^{\pi^{*},p}\geq\Phi(\pi)-\Phi(\pi^{*})\geq 0$, and Assumption 2. By recursively, expanding we get the result. ∎
@@ Appendix D Robust policy evaluator based on KL divergence
8:  $P^{*}=p^{0}(s,a,.)\exp{\left(\alpha\frac{V[.]}{C_{KL}}\right)}$
@@ E.1.2 Discussions of the result
The results highlight RNPG’s ability to consistently learn robust and safe policies while outperforming RPPG and EPIRC_PGS in terms of both reliability and computational efficiency, even under adverse environmental dynamics.
@@ E.3.2 Discussion of results
The results obtained of the given environment is depicted in Figure 6. As seen in Figure 6(b), all the three algorithms successfully learns the feasible policies. However, on observing Figure 6(a), we can clearly notice the dominance of RNPG by learning policies with better rewards. From Table 11, we see that for the frozen lake environment, the computation time of RPPG and RNPG is almost comparable. However, RNPG is atleast 33x as faster as EPIRC_PGS for $\gamma=0.9$ and almost 44x faster than EPIRC_PGS for $\gamma=0.995$.
The key takeaway from this enviornment is to show that even with added obstacles randomly, the agent can find a feasible high objective return policy as compared to RPPG and EPIRC_PGS
@@ E.4.1 Environment description
Figure 7: Comparison of RNPG, RPPG and EPIRC-PGS on Garbage collector environment
@@ E.4.2 Discussion of results
Wall-clock time comparisons (in secs)
Environment Name RPPG RNPG (our) EPIRC-PGS ($\gamma$ vals.)
@@
CRS 81.36 48.574 190.53 228.65 301.21
Garnet 83.36 78.406 390.21 516.23 653.14
Modified Frozenlake 162.35 160.31 453.7 561.47 620.12
Garbage collector 175.25 177.13 344.87 400.13 489.54
Table 11: Comparison of execution times averaged over multiple runs between RNPG, RPPG, and EPIRC-PGS (inner loop T = 100 and outer-loop K=10)
@@ F.2 Robust Projected Policy Gradient (RPPG)
Appendix G escalating to higher space
Algorithm 5 Robust Constrained Actor-Critic (RCAC)
1:Input: Environment $\mathcal{E}$, learning rates $\alpha_{V},\alpha_{C},\alpha_{\pi}$, discount factor $\gamma$, perturbation level $\delta$, constraint bound $b$, trade-off parameter $\lambda$, total episodes NN
2:Initialize actor $\pi_{\theta}$, critic $V_{\psi}$ (for reward), critic $C_{\phi}$ (for constraint)
3:for each training epoch do
4:  for each batch of size $B$ do
5:   Initialize replay buffer $\mathcal{D}\leftarrow\emptyset$
6:   Reset environment and observe initial state $s$
7:   for $i=1$ to $B$ do
8:     Sample action $a\sim\pi_{\theta}(\cdot|s)$
9:     Execute $a$ in $\mathcal{E}$; observe $r$, $c$, $s^{\prime}$, done flag
10:     Store transition $(s,a,\log\pi(a|s),r,c,s^{\prime},done)$ in $\mathcal{D}$
11:     if donedone then reset $s\leftarrow\mathcal{E}.\text{reset}()$
@@
13:     end if
14:   end for
15:   Convert replay buffer $\mathcal{D}$ to tensors
16:   Compute advantage estimates:
$\text{adv}_{r}=r+\gamma(1-\text{done})V_{\psi}(s^{\prime})-V_{\psi}(s)$
$\text{adv}_{c}=c+\gamma(1-\text{done})C_{\phi}(s^{\prime})-C_{\phi}(s)$
17:   Compute perturbed advantages:
[…]
18:   Update critic parameters:
$\psi\leftarrow\psi-\alpha_{V}\nabla_{\psi}\mathcal{L}_{V}\text{ and }\quad\phi\leftarrow\phi-\alpha_{C}\nabla_{\phi}\mathcal{L}_{C}$
where $\mathcal{L}_{V}=\text{MSE}(r+\gamma V(s^{\prime}),V(s))$, $\mathcal{L}_{C}=\text{MSE}(c+\gamma C(s^{\prime}),C(s))$
19:   Select constraint-violating component:
$\text{choice}=\arg\max\left(\frac{V(s)}{\lambda},C(s)-b\right)$
20:   Compute actor loss:
$\mathcal{L}_{\pi}=-\log\pi(a|s)\cdot\delta_{r}\cdot(1-choice)-\log\pi(a|s)\cdot\delta_{c}\cdot(choice),$
21:   Update actor parameters:
$\theta\leftarrow\theta-\alpha_{\pi}\nabla_{\theta}\mathcal{L}_{\pi}$
22:  end for
23:end for

### v3

@@ Appendix A Proof of Proposition 1
$\Phi(\pi)-\Phi(\hat{\pi}^{*})\leq J_{c_{i}}^{\pi,p}-\max_{P}J_{c_{i}}^{\hat{\pi}^{*},P}.$ (29)
We now bound the right-hand side, and assume that $p^{*}=\arg\max J_{c_{i}}^{\hat{\pi}^{*},P}$
$V_{c_{i}}^{\pi}(\rho)-V_{c_{i}}^{\hat{\pi}^{*}}(\rho)=V^{\pi}_{c_{i}}(\rho)-\mathbb{E}_{s\sim\rho}\mathbb{E}_{\hat{\pi}^{*}}[c_{i}(s,a)+\gamma\sum_{s^{\prime}}p^{*}(s^{\prime}|s,a)V^{\pi^{*}}_{c_{i}}(s^{\prime})]$
$=\mathbb{E}_{s\sim\rho}\mathbb{E}_{a\sim\hat{\pi}^{*}}[V_{c_{i}}^{\pi}(\rho)-c_{i}(s,a)+\gamma\sum_{s^{\prime}}p(s^{\prime}|s,a)V^{\pi}_{c_{i}}(s^{\prime})]$
$-\mathbb{E}_{s\sim\rho}\mathbb{E}_{\hat{\pi}^{*}}\gamma[\sum_{s^{\prime}}p^{*}(s^{\prime}|s,a)V^{\hat{\pi}^{*}}_{c_{i}}(s^{\prime})-\sum_{s^{\prime}}p(s^{\prime}|s,a)V^{\pi}_{c_{i}}(s^{\prime})]$
$\leq\sum_{s\sim\rho}\langle Q_{c_{i}}^{\pi},\pi-\hat{\pi}^{*}\rangle-\gamma\mathbb{E}_{s\sim\rho}\mathbb{E}_{\hat{\pi}^{*}}[\sum_{s^{\prime}}p(s^{\prime}|s,a)(V^{\hat{\pi}^{*}}_{c_{i}}(s^{\prime})-V^{\pi}_{c_{i}}(s^{\prime}))]$
where the inequality follows from the fact that $p^{*}=\arg\max\sum_{s^{\prime}}p^{*}(s^{\prime}|s,a)V_{c_{i}}^{\hat{\pi}^{*}}(s^{\prime})$. Hence,
$V_{c_{i}}^{\pi}(\rho)-V_{c_{i}}^{\hat{\pi}^{*}}(\rho)\leq\sum_{s\sim\rho}\langle Q_{c_{i}}^{\pi},\pi-\hat{\pi}^{*}\rangle+\beta\mathbb{E}_{s\sim\rho}\mathbb{E}_{\hat{\pi}^{*}}[\sum_{s^{\prime}}p_{0}(s^{\prime}|s,a)(V^{\pi}_{c_{i}}(s^{\prime})-V^{\hat{\pi}^{*}}_{c_{i}}(s^{\prime}))].$ (30)
where we use the fact that $V_{c_{i}}^{\pi}(s^{\prime})\geq V_{c_{i}}^{\hat{\pi}^{*}}(s^{\prime})$ by Assumption 3, and Assumption 2. By recursively, expanding we get the result. ∎
@@ Appendix D Robust policy evaluator based on KL divergence
8:  $P^{*}=\frac{p^{0}[s,a,.]\exp{\left(\frac{V[.]}{C_{KL}}\right)}}{\sum_{s^{{}^{\prime}}\in\mathcal{S}}p^{0}[s,a,s^{{}^{\prime}}]\exp{\left(\frac{V[s^{{}^{\prime}}]}{C_{KL}}\right)}}~\forall(s,a)$
@@ E.1.2 Discussions of the result
The results highlight RNPG’s ability to consistently learn robust and safe policies while outperforming RPPG and EPIRC-PGS in terms of both reliability and computational efficiency, even under adverse environmental dynamics.
@@ E.3.2 Discussion of results
The results obtained of the given environment is depicted in Figure 6. As seen in Figure 6(b), all the three algorithms successfully learns the feasible policies. However, on observing Figure 6(a), we can clearly notice the dominance of RNPG by learning policies with better rewards. From Table 11, we see that for the frozen lake environment, the computation time of RPPG and RNPG is almost comparable. However, RNPG is atleast 33x as faster as EPIRC-PGS for $\gamma=0.9$ and almost 44x faster than EPIRC-PGS for $\gamma=0.995$.
The key takeaway from this enviornment is to show that even with added obstacles randomly, the agent can find a feasible high objective return policy as compared to RPPG and EPIRC-PGS
@@ E.4.1 Environment description
Figure 7: Comparison of RNPG, RPPG and EPIRC-PGS (or, EPIRC_PGS) on Garbage collector environment
@@ E.4.2 Discussion of results
Best policy’s objective and constraint return comparison
Env Nm RPPG RNPG EPIRC-PGS ($\gamma$ vals.)
@@
$v_{f}$ $c_{f}$ $v_{f}$ $c_{f}$ $v_{f}$ $c_{f}$ $v_{f}$ $c_{f}$ $v_{f}$ $c_{f}$
CRS, $b_{1}=42.5$ 120.3 43.1 102.1 42.2 127.2 44.8 121.2 48.1 123.4 46.4
Garnet, $b_{1}=-90$ 102.6 -96.5 208.1 -98.7 100.3 -112.4 100.6 -108.3 102.4 -110.2
MFL, $b_{1}=52.5$ 61.4 53.1 109.2 39.2 62.3 53.7 60.3 57.7 66.3 51.7
GC, $b_{1}=52.5$ 161.2 49.2 172.2 22.1 131.2 52.0 142.5 55.6 138.1 54.1
Table 11: Comparison of the best policy objective ($v_{f}$) and constraint function ($c_{f}$) values. $b_{1}$ indicates the threshold value. RNPG not only achieves the best value, but also gives a feasible policy.
@@ F.2 Robust Projected Policy Gradient (RPPG)
Appendix G Extension to Continuous state space (Robust Constrained Actor Critic)
We present our robust constrained actor–critic framework designed for the function approximation setting as discussed in Section 6.1. The model comprises two critic networks—one estimating the reward value function $J_{c_{0}}$ and the other corresponding to the Constraint value function $J_{c_{1}}$. Although we focus on a single constraint for clarity, the framework readily generalizes to handle multiple constraints. In addition to the critic networks, an actor network is employed to generate actions based on the current state. To model distributional robustness, we consider IPM as described in Section 6.1.
Let us consider the critic network be parameterized by $w$ where each layer contains $d$ paramters including the bias (i.e, $w^{l}_{1:d}$ where $w^{l}_{1}$ is assumed to be the bias term in the $l$-th layer with $l\in\{1,2,\ldots,L\}$). Overall approach is depicted in Algorithm 5.
Algorithm 5 Robust Constrained Actor Critic (RCAC)
1:Input: $T,\rho,b$
2:Initialization: $w_{r}$ (for objective estimation), $w_{c}$ (for constraint estimation) and $\theta^{0}$ for actor network parameterization
3:for $t=0,\ldots,T-1$ do
4:  Get estimate for $J_{r}=\langle\rho,V_{w_{r}}(s)\rangle$ and $J_{c}=\langle\rho,V_{w_{c}}(s)\rangle$
5:  $ch=\arg\max(J_{r}/\lambda,(J_{c}-b))$
@@ Appendix G Extension to Continuous state space (Robust Constrained Actor Critic)
7:  Update $\theta$ using $\theta^{t}=\theta^{t-1}+\alpha.\mathbb{E}\left[\nabla_{\theta}\log(\pi_{\theta}(a|s)).(Q_{ch}(s,a)-V_{ch}(s))\right]$ (We change this step to Natural Policy Gradient update for RCAC_NPG).
8:end for
At each step, we get the Value function estimate $V_{r}$ and $V_{c}$ from the respective Critic networks. After obtaining both, we make a choice as to whether to update the constraint critic parameters or the objective critic parameters. For the selected critic, we find the target value function using the robust bellman operator along with a guided regularization term on the last layer only [54]. We compute the robust value according to (15).
For our experiments, we chose the famous Cartpole environment, where the intial state is fixed and deterministic so $\rho(i)$ is a unit vector.(However, it can be extended to different distribution.)
$\rho(s)=\begin{cases}1&\text{if $s=s_{0}$}\\ 0&\text{otherwise}\end{cases}$ (47)
In our study, we introduced uncertainty in the next-state transition after each action. While alternative sources of uncertainty could be incorporated—such as perturbing the executed actions or simulating external disturbances (e.g., wind forces acting on the cart)—we focused on state transition perturbations because they have a more direct and analyzable impact on value estimation. Perturbing the action space was deemed less meaningful in this environment, as the action set is discrete with only two possible values, making the resulting learning challenge comparatively trivial. The detailed results and observations are presented in the following subsection.
G.1 Results and discussion
In this sub-section, we list the results obtained when we tried our algorithm against the standard cartpole-v1 environment available in gymnasium library. The cartpole-v1 algorithm comprises of a continuous state space having 4 components and two discrete actions. We introduce uncertainity by adding noise to the next state obtained after taking an action. The noise is adding a uniform value between 0 and 0.1 to the original next state value ($s^{{}^{\prime}}=s^{{}^{\prime}}+Unif([0,0.1])$) and then clipped it between the predefined bounds of cartpole environment.
We divided the experiments results into two phases. The first is the training phase (depicted by Figures 8) and the second is during the testing phase (depicted by Figures 9). During the training phase, we only train the robust variants RCAC, RCAC with NPG, Robust CRPO, EPIRC-PGS by considering $\delta=0.04$. However, for constrained actor-critic (CAC), we do not train the robust version. During the training phase (Figure 8), apart from EPIRC-PGS, all the algorithms perform similarly in terms of reward and the cost value function (We highlight our two algorithms RCAC with NPG, and RCÅC, separately in figure 10). EPIRC-PGS did not converge and could not complete the entire episode highlighting that binary search approach is not possible to scale for large state-space. However, when these algorithms were tested (Figure 9) on the environment having a perturbation uniformly between 0 and 0.04, the performance of CAC is unstable and did not provide any feasible policy. Only, our proposed approaches achieve feasibility while being close to the optimality. Robust CRPO also violates the constraint (slightly) while achieving less reward compared to our approaches. The performances of the algorithms is compactly represented in table 12.
It is also important to note down the wall clock time for the various algorithms. EPIRC_PGS takes the highest wall clock time approximately 24029.013 seconds which is nearly $4\times$ of the time taken for the other algorithms namely RCAC with NPG (7175.31 seconds), CAC(6815.89 seconds), RCAC (5975.65 seconds) and Robust CRPO (4275.67 seconds) in decreasing order of the wall clock time requirements. 9.
(a) Value function comparison
(b) Cost function comparison
Figure 8: Comparison of RCAC, RCAC_NPG, robust CRPO and other standard Constrained MDP solutions on the Cartpole problem during training phase
(a) Value function comparison
(b) Cost function comparison
Figure 9: Comparison between RCAC, CRPO and Vanilla Constrained Actor Critic during testing period $\delta=0.04$ (deflection from the nominal model)
Algorithm Average reward value Average cost value
Constrained Actor Critic 495.23 294.6248
RCAC 494.81 177.8432
RCAC with NPG 495.0 197.9937
Robust CRPO 488.06930 213.473
EPIRC_PGS 114.8 53.79
Table 12: Tabular comparisons of the average value function and cost function during the testing phase. CAC although returns policies with high objective function but the actions are unsafe as can be inferred from the high constraint function (safety baseline is 200)
(a) Value function comparison
(b) Cost function comparison
Figure 10: The comparison plots between our two main variants of Robust Constrained Actor Critic variants (RCAC and RCAC_NPG)